OpenStudy (anonymous):

simplify the radical expressions rad72- rad50

OpenStudy (saifoo.khan):

\[\sqrt{72} \to 6 \sqrt2\] \[\sqrt{50} \to 5 \sqrt2\]

OpenStudy (saifoo.khan):

Can you solve now?

OpenStudy (anonymous):

I'm sorry I am completely lost when it comes to radicals

OpenStudy (lgbasallote):

i'll just show a more detailed thingy of what he did \[\sqrt{72} \rightarrow \sqrt{(36)(2)} \rightarrow 6\sqrt 2\] \[\sqrt{50} \rightarrow \sqrt{(25)(2)} \rightarrow 5\sqrt 2\]

OpenStudy (anonymous):

so